# Towards Galactic Archaeology with Inferred Ages of Giant Stars From Gaia Spectra

### v1

@@ Abstract
In the era of Gaia, the accurate determination of stellar ages is transforming Galactic archaeology. We demonstrate the feasibility of inferring stellar ages from Gaia’s RVS spectra and the BP/RP (XP) spectrophotometric data, specifically for red giant branch and high-mass red clump stars. We successfully train two machine learning models, dubbed SIDRA: Stellar age Inference Derived from Gaia spectRA to predict the age. The SIDRA-RVS model uses the RVS spectra and SIDRA-XP the stellar parameters obtained from the XP spectra by Fallows & Sanders (2024). Both models use BINGO, an APOGEE-derived stellar age from Ciucă et al. (2021) as the training data. SIDRA-RVS estimates ages of stars whose age is around $\tau_{\mathrm{BINGO}}=10$ Gyr with a standard deviation of residuals of $\sim$ 0.12 dex in the unseen test dataset, while SIDRA-XP achieves higher precision with residuals $\sim$ 0.064 dex for stars around $\tau_{\mathrm{BINGO}}=10$  Gyr. Since SIDRA-XP outperforms SIDRA-RVS, we apply SIDRA-XP to analyse the ages for 2,218,154 stars. This allowed us to map the chronological and chemical properties of Galactic disc stars, revealing distinct features such as the Gaia-Sausage-Enceladus merger and a potential gas-rich interaction event linked to the first infall of the Sagittarius dwarf galaxy. This study demonstrates that machine learning techniques applied to Gaia’s spectra can provide valuable age information, particularly for giant stars, thereby enhancing our understanding of the Milky Way’s formation and evolution.
@@
pubyear: 2015pagerange: Towards Galactic Archaeology with Inferred Ages of Giant Stars From Gaia SpectraB
@@ 1 Introduction
With the advent of large spectroscopic surveys, such as the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Majewski et al., 2017), Galactic Archaeology with HERMES (GALAH, De Silva et al., 2015), Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST, Cui et al., 2012) and the precise astrometry from the Gaia mission (Gaia Collaboration et al., 2016, 2023a), precise determination of stellar chemical abundances, motions and estimated ages—known as chrono-chemodynamical information—is revolutionising Galactic archaeology. It marks an initial stride in empirically comprehending the formation and evolution of our Galaxy that have led to its current structure.
@@
An alternative method (but also model-dependent) for determining the ages of stars, including red giants, is through asteroseismology (Chaplin & Miglio, 2013). Asteroseismology entails directly measuring the seismic frequencies of a star to derive accurate measurements of its internal properties and structure, including radius and mass (see Aerts, 2021, for a review). Since the mass of the star determines its main sequence lifetime, the mass of a low-mass red giant can effectively reveal its age precisely (e.g., see Miglio, 2012). Asteroseismic measurements of the mass and ages of the stars, including giants, are conducted with Kepler (Koch et al., 2010; Gilliland et al., 2010; Bedding et al., 2010; Ceillier et al., 2017; Yu et al., 2018), Convection, Rotation, and Planetary Transits (CoRoT, Auvergne et al., 2009; Miglio et al., 2009; Mosser et al., 2010; Kallinger et al., 2010; Valentini et al., 2016), K2 (Howell et al., 2014; Zinn et al., 2022) and the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014; Mackereth et al., 2021; Hon et al., 2021). However, asteroseismology requires a significant amount of precise time-series photometry and computational resources, making it feasible only for a limited number of stars at a time.
Masseron & Gilmore (2015) demonstrated that photospheric carbon and nitrogen abundances (specifically [C/N]) are a good indicator of the mass of giant stars, thereby providing the possibility to estimate their age from stellar spectra. This finding enables the use of extensive spectroscopic survey data, either directly from the spectra or through derived spectroscopic-parameters, to predict stellar ages. However, this method requires calibration with precisely measured ages or masses by asteroseismology. Building on this, Martig et al. (2016) showed that it is possible to infer ages from the well-calibrated APOGEE DR12 spectroscopic derived parameters. They demonstrated that the masses of red-giant stars can be precisely predicted based on their [C/N] abundances, along with their spectroscopic stellar labels, including $T_{\mathrm{eff}}$, log g and [Fe/H] (see also Mackereth et al., 2017; Ciucă et al., 2021; Anders et al., 2023). Similarly, Ness et al. (2016) found that CNO molecular lines, specifically CN and CO, within the APOGEE survey also encompass mass-related and, consequently, age-related information and used the spectra directly to predict ages. Extending this approach to other surveys, Ho et al. (2017) illustrated that the CH and CN characteristics in the blue region of the LAMOST spectra are indicative of mass, and hence age (see also He et al., 2022).
@@
Presently, Gaia’s third data release (DR3, Gaia Collaboration et al., 2023a) is producing mean Radial Velocity Spectrometer (RVS) spectra for 1 million well-behaved objects, with a significant portion expected to be red giants. These are all-sky high resolution spectra (approximately R = 11,500), enabling a comprehensive study of Milky Way archaeology, structure and evolution (Gaia Collaboration et al., 2023c). Chemical abundances and atmospheric stellar parameters have been estimated from the RVS spectra by the Gaia collaboration (Recio-Blanco et al., 2023).
Additionally, low-resolution spectrophotometric measurements, known as Gaia BP/RP spectra (hereafter XP spectra), are available for over about 220 million objects as part of the third data release of Gaia DR3. These spectra cover the optical wavelength range from 330 nm to 1050 nm with a wavelength-dependent resolution between 30 and 100 (Carrasco et al., 2021; De Angeli et al., 2023; Gaia Collaboration et al., 2023b). Despite their low resolution, these spectra provide sufficient information to accurately estimate fundamental stellar parameters, such as effective temperature, surface gravity and metallicity (Liu et al., 2012; Witten et al., 2022; Andrae et al., 2023; Martin et al., 2023), with the added potential to derive more detailed abundance measurements, including elements like carbon and alpha elements (Gavel et al., 2021; Witten et al., 2022; Lucey et al., 2023; Sanders & Matsunaga, 2023; Li et al., 2024). Further Guiglion et al. (2024) used machine learning models to derive the stellar parameters more precisely, using the combined information of the RVS spectra, photometry, parallax and the XP data.
Building on previous studies that extracted detailed abundance information from XP spectra (e.g. Witten et al., 2022; Li et al., 2024), Fallows & Sanders (2024, hereafter FS24) developed a feed-forward neural network to predict stellar parameters and abundances using XP spectra, along with photometry from Gaia, the Two Micron All Sky Survey (2MASS, Skrutskie et al., 2006) and Wide-field Infrared Survey Explorer (unWISE, Schlafly et al., 2019). Their model carefully considers uncertainties in both inputs and outputs, as well as additional model uncertainty. They have published their estimates of $T_{\mathrm{eff}}$, log $g$ and various abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and [$\mathrm{\alpha/M}$].
@@ 2 BINGO giants sample
Ciucă et al. (2024) estimated the age for red giant branch (RGB) and high-mass red-clump (RC) stars using BINGO as described in Ciucă et al. (2021). BINGO is a Bayesian Neural Network model designed to map APOGEE stellar parameters, such as $T_{\mathrm{eff}}$, log g, [Fe/H], [Mg/Fe], [C/Fe] and [N/Fe], to stellar age. The model is trained using APOGEE DR17 data in conjunction with Kepler asteroseismology data, with stellar parameters from APOGEE DR17 as input features and asteroseismic age data from Miglio et al. (2021) as the output, while accounting for the observational uncertainties during the training process. Ciucă et al. (2024) limited the training data to the RGB stars and high-mass ($>1.8~{}\mathrm{M}_{\odot}$) RC stars, where asteroseismology can measure the initial mass of the stars confidently, because of their negligible mass loss.
@@
We cross-match the APOGEE BINGO stars with Gaia DR3 stars with RVS spectra, and obtain 8,859 stars with RVS spectra and high-precision BINGO ages (BINGO-RVS data hereafter). Their distribution in the Kiel diagram is seen in the left panel of Fig. 1. The high-mass RC stars (younger stars) are focused mainly around 2.7 $<\log$ $g<$ 3, 3.69 $<\mathrm{log}$ $T_{\mathrm{eff}}<$ 3.71 with ages between $-0.5<\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}\mathrm{[Gyr]})<0.25$. This trend is typical among younger RC stars, because those in the core helium burning phase tend to have higher $\log~{}g$ compared to the lower mass RC stars.
We also construct our ground-truth BINGO giant sample to build and evaluate a machine-learning model to estimate the stellar age for the giant stars from Gaia’s XP stellar parameters derived by FS24. The selected dataset from FS24 XP stellar parameter data consists of stars that meet specific accuracy criteria and parameter ranges. We select only the stars whose stellar parameters uncertainties are within their reported median formal uncertainties, i.e. an effective temperature uncertainty, $T_{\mathrm{eff}_{\mathrm{unc}}}<69.1$ K, surface gravity uncertainty, log $g_{\mathrm{unc}}<0.14$ dex, and metallicity uncertainties, $[\mathrm{Fe}/\mathrm{H}]_{\mathrm{unc}}<0.068$ dex, and $[\mathrm{\alpha}/\mathrm{M}]_{\mathrm{unc}}<0.029$ dex. In cases where duplicate entries exist, the star with the smaller $T_{\mathrm{eff}_{\mathrm{unc}}}$ is selected. We then cross-match these data with the APOGEE stars with the BINGO ages from Ciucă et al. (2021), and construct our ground-truth data (a total of 12,500 stars, BINGO-FS24 data hereafter). In this work, we use the FS24 data with only the RVS radial velocity data, as we find the predicted stellar parameters to be more reliable when restricted to this dataset. Additionally, this dataset offers a similar magnitude range to our cross-matched RVS spectra sample, enabling a fair comparison.
@@ 3.1 SIDRA-RVS’s methodology
We develop SIDRA-RVS to predict the ages from the Gaia spectra of RGB and high-mass RC stars. For the analysis of RVS spectra, we use an eXtreme Gradient Boosting (XGBoost, Chen & Guestrin, 2016), a tree-based machine-learning algorithm that constructs a sequence of decision trees to provide predictions for regression tasks. Each decision tree corrects the errors of the previous ones, contributing to improving the overall model’s accuracy. The key abilities and features of this model make it well matched to accurately predict stellar ages. For example, XGBoost exhibits excellent performance in managing extensive data sets, demonstrating notably faster processing speeds than ANNs. It is a flexible algorithm for dealing with missing data instances, making it suitable for the characteristics of our RVS data that sometimes include missing data points in the spectra. RVS spectra are observed within the 846-870 nm filter in the observed wavelength range. Before making the combined spectrum, they are Doppler-shifted to the rest frame. Because Gaia DR3 provides the data in a fixed wavelength range in the rest frame, and stars with a large radial velocity have a substantial shift in the wavelength range, one end of the spectrum lead to no data for such stars. There are sometimes a lack of the data in the middle of the spectrum, when all the observed data in that wavelength bin have been corrupted for various reasons, e.g. due to defective pixel. Moreover, XGBoost is an ensemble method which combines the predictions of multiple weak regression models or learners known as ‘trees’ to create a more robust and accurate model. This ensemble strategy improves predictive performance and mitigates the risk of overfitting. This was shown by Borisov et al. (2022) who benchmarked various regression algorithms for heterogeneous data sets and found that algorithms based on gradient-boosted tree ensembles perform better than deep-learning models in supervised learning tasks. This model has been utilised for similar previous research, including the work of Anders et al. (2023), who employed the XGBoost algorithm to assess the age and chemical abundances of APOGEE red-giant stars.
@@ 3.2 Age inference from RVS spectra using SIDRA-RVS
Figure 2: SIDRA-RVS age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the training set colour-coded by [M/H] (left panel) and [$\mathrm{\alpha/M}$] (right panel) obtained from APOGEE. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
@@
In this section, we explore the performance of the stellar-age measurements with the RVS spectra from our sample of RGB and high-mass RC stars using XGBoost. Fig. 2 illustrates the comparison between the ground-truth BINGO ages, $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$, and the corresponding SIDRA-RVS predicted ages, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, for the stars in the training data, while Fig. 3 shows the same results for the test data. We adopt $80\%$ of our cross-matched BINGO-RVS data for the training data, and the rest of the data are used as the unseen test data after the model is trained. Note that the data augmentation outlined in Sec. 3.1 is applied exclusively to the training data. As a result, Fig. 2 displays a larger and more uniformly distributed dataset with respect to age. In contrast, the test data shown in Fig. 3 illustrates the imbalance in the data. In both the training and testing sets, despite some variations in the predictions, the majority of data points closely follow the identity line. This agreement indicates that SIDRA-RVS effectively predicts the ages of the selected giant stars only from the RVS spectra. The bottom panel represents the residuals between SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and the ground truth BINGO $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$. In the case of the test data shown in Fig. 3, the dispersion of residuals is significantly higher for stars with lower ages ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})<0.2$) due to the scarcity of young stars in the test dataset, which leads to a dominance of Poisson noise. In contrast, the residual dispersion for the training data in Fig. 2 remains nearly constant across all ages. It is important to note that without the data augmentation applied to the training data, a severe overestimation of the age of young stars is observed in both the training and test data, as there is a lack of young stars ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})<0.2$). While the mean deviation for the test data still indicates an overestimation of age for young stars, the performance is significantly improved compared to a model trained without data augmentation. As a result , the mean of the standard deviation of residual is 0.10 dex for stars ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})>0.2$) in the testing set and there are no strong systematic offsets. The standard deviation of the residuals is around 0.09 dex for the training data and 0.12 dex for the test data around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})=1$. This means that the model prediction for the unseen data is as good as the prediction for the training data.
@@
Figure 3: SIDRA-RVS age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the testing set colour-coded by [M/H] (left panel) and [$\mathrm{\alpha/M}$] (right panel) obtained from APOGEE. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
The colours of the dots in Figs. 2 and 3 indicate [M/H] (left panel) and [$\mathrm{\alpha/M}$] (right panel). In general, the older age stars have lower [M/H] and higher $[\alpha/\mathrm{M}]$, and the BINGO ages show the clear correlation with [M/H] and $[\alpha/\mathrm{M}]$, as shown in Ciucă et al. (2021). The SIDRA ages also follow the trend. However, it is not as strong as the trend seen for the BINGO ages, and the stars with similar SIDRA age have a wider variety of [M/H] and [$\alpha$/M]. This ensures that SIDRA is not solely learning the [$\mathrm{\alpha/M}$]-age or [M/H]-age trends, as these trends are more pronounced for BINGO ages.

### v2

@@ Abstract
In the era of Gaia, the accurate determination of stellar ages is transforming Galactic archaeology. We demonstrate the feasibility of inferring stellar ages from Gaia’s RVS spectra and the BP/RP (XP) spectrophotometric data, specifically for red giant branch and high-mass red clump stars. We successfully train two machine learning models, dubbed SIDRA: Stellar age Inference Derived from Gaia spectRA to predict the age. The SIDRA-RVS model uses the RVS spectra and SIDRA-XP the stellar parameters obtained from the XP spectra. Both models use BINGO, an APOGEE-derived stellar age as the training data. SIDRA-RVS estimates ages of stars whose age is around $\tau_{\mathrm{BINGO}}=10$ Gyr with a standard deviation of residuals of $\sim$ 0.12 dex in the unseen test dataset, while SIDRA-XP achieves higher precision with residuals $\sim$ 0.064 dex for stars around $\tau_{\mathrm{BINGO}}=10$  Gyr. Since SIDRA-XP outperforms SIDRA-RVS, we apply SIDRA-XP to analyse the ages for 2,218,154 stars. This allowed us to map the chronological and chemical properties of Galactic disc stars, reproducing the known distinct features such as the Gaia-Sausage-Enceladus merger and a potential gas-rich interaction event linked to the first infall of the Sagittarius dwarf galaxy. This study demonstrates that machine learning techniques applied to Gaia’s spectra can provide valuable individual age information, particularly for giant stars, thereby enhancing our understanding of the Milky Way’s formation and evolution.
@@
pubyear: 2015pagerange: Towards Galactic Archaeology with Inferred Ages of Giant Stars From Gaia SpectraC
@@ 1 Introduction
With the advent of large spectroscopic surveys, such as the Apache Point Observatory Galactic Evolution Experiment (APOGEE, Majewski et al., 2017), Galactic Archaeology with HERMES (GALAH, De Silva et al., 2015), Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST, Cui et al., 2012) and the precise astrometry from the Gaia mission (Gaia Collaboration et al., 2016, 2023c), precise determination of stellar chemical abundances, motions and estimated ages—known as chrono-chemodynamical information—is revolutionising Galactic archaeology. It marks an initial stride in empirically comprehending the formation and evolution of our Galaxy that have led to its current structure.
@@
An alternative method (but also model-dependent) for determining the ages of stars, including red giants, is through asteroseismology (Chaplin and Miglio, 2013). Asteroseismology entails directly measuring the seismic frequencies of a star to derive accurate measurements of its internal properties and structure, including radius and mass (see Aerts, 2021, for a review). Since the mass of the star determines its main sequence lifetime, the mass of a low-mass red giant can effectively reveal its age precisely (e.g., see Miglio, 2012). Asteroseismic measurements of the mass and ages of the stars, including giants, are conducted with Kepler (Koch et al., 2010; Gilliland et al., 2010; Bedding et al., 2010; Ceillier et al., 2017; Yu et al., 2018), Convection, Rotation, and Planetary Transits (CoRoT, Auvergne et al., 2009; Miglio et al., 2009; Mosser et al., 2010; Kallinger et al., 2010; Valentini et al., 2016), K2 (Howell et al., 2014; Zinn et al., 2022) and the Transiting Exoplanet Survey Satellite (TESS, Ricker et al., 2014; Mackereth et al., 2021; Hon et al., 2021). However, asteroseismology requires a significant amount of precise time-series photometry and computational resources, making it feasible only for a limited number of stars at a time.
Masseron and Gilmore (2015) demonstrated that photospheric carbon and nitrogen abundances (specifically [C/N]) are a good indicator of the mass of giant stars, thereby providing the possibility to estimate their age from stellar spectra. This finding enables the use of extensive spectroscopic survey data, either directly from the spectra or through derived spectroscopic-parameters, to predict stellar ages. However, this method requires calibration with precisely measured ages or masses by asteroseismology. Building on this, Martig et al. (2016) showed that it is possible to infer ages from the well-calibrated APOGEE DR12 spectroscopic derived parameters. They demonstrated that the masses of red-giant stars can be precisely predicted based on their [C/N] abundances, along with their spectroscopic stellar labels, including $T_{\mathrm{eff}}$, log g and [Fe/H] (see also Mackereth et al., 2017; Ciucă et al., 2021; Anders et al., 2023). Similarly, Ness et al. (2016) found that CNO molecular lines, specifically CN and CO, within the APOGEE survey also encompass mass-related and, consequently, age-related information and used the spectra directly to predict ages. Extending this approach to other surveys, Ho et al. (2017) illustrated that the CH and CN characteristics in the blue region of the LAMOST spectra are indicative of mass, and hence age (see also He et al., 2022).
@@
Presently, Gaia’s third data release (DR3, Gaia Collaboration et al., 2023c) is producing mean Radial Velocity Spectrometer (RVS) spectra for 1 million well-behaved objects, with a significant portion expected to be red giants. These are all-sky high resolution spectra (approximately R = 11,500), enabling a comprehensive study of Milky Way archaeology, structure and evolution (Gaia Collaboration et al., 2023b). Chemical abundances and atmospheric stellar parameters have been estimated from the RVS spectra by the Gaia collaboration (Recio-Blanco et al., 2023).
Additionally, low-resolution spectrophotometric measurements, known as Gaia BP/RP spectra (hereafter XP spectra), are available for over about 220 million objects as part of the third data release of Gaia DR3. These spectra cover the optical wavelength range from 330 nm to 1050 nm with a wavelength-dependent resolution between 30 and 100 (Carrasco et al., 2021; De Angeli et al., 2023; Gaia Collaboration et al., 2023a). Despite their low resolution, these spectra provide sufficient information to accurately estimate fundamental stellar parameters, such as effective temperature, surface gravity and metallicity (Liu et al., 2012; Witten et al., 2022; Andrae et al., 2023; Martin et al., 2024), with the added potential to derive more detailed abundance measurements, including elements like carbon and alpha elements (Gavel et al., 2021; Witten et al., 2022; Lucey et al., 2023; Sanders and Matsunaga, 2023; Li et al., 2024). Further Guiglion et al. (2024) used machine learning models to derive the stellar parameters more precisely, using the combined information of the RVS spectra, photometry, parallax and the XP data.
Building on previous studies that extracted detailed abundance information from XP spectra (e.g. Witten et al., 2022; Li et al., 2024), Fallows and Sanders (2024, hereafter FS24) developed a feed-forward neural network to predict stellar parameters and abundances using XP spectra, along with photometry from Gaia, the Two Micron All Sky Survey (2MASS, Skrutskie et al., 2006) and Wide-field Infrared Survey Explorer (unWISE, Schlafly et al., 2019). Their model carefully considers uncertainties in both inputs and outputs, as well as additional model uncertainty. They have published their estimates of $T_{\mathrm{eff}}$, log $g$ and various abundance ratios, such as [C/Fe], [Fe/H], [N/Fe] and $\mathrm{\alpha/M}$].
@@ 2 BINGO giants sample
Ciucă et al. (2024) estimated the age for red giant branch (RGB) and high-mass red-clump (RC) stars using BINGO as described in Ciucă et al. (2021). BINGO is a Bayesian Neural Network model designed to map APOGEE stellar parameters, such as $T_{\mathrm{eff}}$, log g, [Fe/H], [Mg/Fe], [C/Fe] and [N/Fe], to stellar age. The model is trained using APOGEE DR17 data in conjunction with Kepler asteroseismology data, with stellar parameters from APOGEE DR17 as input features and asteroseismic age data from Miglio et al. (2021) as the output, while accounting for the observational uncertainties during the training process. Ciucă et al. (2024) limited the training data to the RGB stars and high-mass ($1.8~\mathrm{M}_{\odot}$) RC stars, where asteroseismology can measure the initial mass of the stars confidently, because of their negligible mass loss.
@@
We cross-match the APOGEE BINGO stars with Gaia DR3 stars with RVS spectra, and obtain 8,859 stars with RVS spectra and high-precision BINGO ages (BINGO-RVS data hereafter). Their distribution in the Kiel diagram is seen in the left panel of Fig. 1. The high-mass RC stars (younger stars) are focused mainly around 2.7 $\log$ $g<$ 3, 3.69 $\mathrm{log}$ $T_{\mathrm{eff}}<$ 3.71 with ages between $0.5<\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}\mathrm{[Gyr]})<0.25$. This trend is typical among younger RC stars, because those in the core helium burning phase tend to have higher $\log~g$ compared to the lower mass RC stars.
We also construct our ground-truth BINGO giant sample to build and evaluate a machine-learning model to estimate the stellar age for the giant stars from Gaia’s XP stellar parameters derived by FS24. The selected dataset from FS24 XP stellar parameter data consists of stars that meet specific accuracy criteria and parameter ranges. We select only the stars whose stellar parameters uncertainties are within their reported median formal uncertainties, i.e. an effective temperature uncertainty, $T_{\mathrm{eff}_{\mathrm{unc}}}<69.1$ K, surface gravity uncertainty, log $g_{\mathrm{unc}}<0.14$ dex, and metallicity uncertainties, $\mathrm{Fe}/\mathrm{H}]_{\mathrm{unc}}<0.068$ dex, and $\mathrm{\alpha}/\mathrm{M}]_{\mathrm{unc}}<0.029$ dex. In cases where duplicate entries exist, the star with the smaller $T_{\mathrm{eff}_{\mathrm{unc}}}$ is selected. We then cross-match these data with the APOGEE stars with the BINGO ages from Ciucă et al. (2021), and construct our ground-truth data (a total of 12,500 stars, BINGO-FS24 data hereafter). In this work, we use the FS24 data with only the RVS radial velocity data, as we find the predicted stellar parameters to be more reliable when restricted to this dataset. Additionally, this dataset offers a similar magnitude range to our cross-matched RVS spectra sample, enabling a fair comparison.
@@ 3.1 SIDRA-RVS’s methodology
We develop SIDRA-RVS to predict the ages from the Gaia spectra of RGB and high-mass RC stars. For the analysis of RVS spectra, we use an eXtreme Gradient Boosting (XGBoost, Chen and Guestrin, 2016), a tree-based machine-learning algorithm that constructs a sequence of decision trees to provide predictions for regression tasks. Each decision tree corrects the errors of the previous ones, contributing to improving the overall model’s accuracy. The key abilities and features of this model make it well matched to accurately predict stellar ages. For example, XGBoost exhibits excellent performance in managing extensive data sets, demonstrating notably faster processing speeds than ANNs. It is a flexible algorithm for dealing with missing data instances, making it suitable for the characteristics of our RVS data that sometimes include missing data points in the spectra. RVS spectra are observed within the 846-870 nm filter in the observed wavelength range. Before making the combined spectrum, they are Doppler-shifted to the rest frame. Because Gaia DR3 provides the data in a fixed wavelength range in the rest frame, and stars with a large radial velocity have a substantial shift in the wavelength range, one end of the spectrum lead to no data for such stars. There are sometimes a lack of the data in the middle of the spectrum, when all the observed data in that wavelength bin have been corrupted for various reasons, e.g. due to defective pixel. Moreover, XGBoost is an ensemble method which combines the predictions of multiple weak regression models or learners known as ‘trees’ to create a more robust and accurate model. This ensemble strategy improves predictive performance and mitigates the risk of overfitting. This was shown by Borisov et al. (2022) who benchmarked various regression algorithms for heterogeneous data sets and found that algorithms based on gradient-boosted tree ensembles perform better than deep-learning models in supervised learning tasks. This model has been utilised for similar previous research, including the work of Anders et al. (2023), who employed the XGBoost algorithm to assess the age and chemical abundances of APOGEE red-giant stars.
@@ 3.2 Age inference from RVS spectra using SIDRA-RVS
Figure 2: SIDRA-RVS age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the training set colour-coded by [M/H] (left panel) and $\mathrm{\alpha/M}$] (right panel) obtained from APOGEE. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
@@
In this section, we explore the performance of the stellar-age measurements with the RVS spectra from our sample of RGB and high-mass RC stars using XGBoost. Fig. 2 illustrates the comparison between the ground-truth BINGO ages, $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$, and the corresponding SIDRA-RVS predicted ages, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$, for the stars in the training data, while Fig. 3 shows the same results for the test data. We adopt $80\%$ of our cross-matched BINGO-RVS data for the training data, and the rest of the data are used as the unseen test data after the model is trained. Note that the data augmentation outlined in Sec. 3.1 is applied exclusively to the training data. As a result, Fig. 2 displays a larger and more uniformly distributed dataset with respect to age. In contrast, the test data shown in Fig. 3 illustrates the imbalance in the data. In both the training and testing sets, despite some variations in the predictions, the predictions generally follow the identity line, though a mild systematic trend is present. Using the binned points in the lower panel of Fig. 3, the residuals show a negative slope of approximately $\text{slope}\approx-0.27\pm 0.01$, corresponding to a change of roughly 0.1 dex in the residuals from the youngest to the oldest stars, indicating that the SIDRA-RVS ages are slightly compressed relative to the BINGO ages. The bottom panel represents the residuals between SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and the ground truth BINGO $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$. In the case of the test data shown in Fig. 3, the dispersion of residuals is significantly higher for stars with lower ages ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})<0.2$) due to the scarcity of young stars in the test dataset, which leads to a dominance of Poisson noise. In contrast, the residual dispersion for the training data in Fig. 2 remains nearly constant across all ages. It is important to note that without the data augmentation applied to the training data, a severe overestimation of the age of young stars is observed in both the training and test data, as there is a lack of young stars ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})<0.2$). While the mean deviation for the test data still indicates an overestimation of age for young stars, the performance is significantly improved compared to a model trained without data augmentation. As a result , the mean of the standard deviation of residual is 0.10 dex for stars ($\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})>0.2$) in the testing set and there are no strong systematic offsets. The standard deviation of the residuals is around 0.09 dex for the training data and 0.12 dex for the test data around $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})=1$. This suggests that the model performs similarly on the test set, although a modest increase in scatter remains.
@@
Figure 3: SIDRA-RVS age predictions, $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ versus the target BINGO age estimation in $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ for the testing set colour-coded by [M/H] (left panel) and $\mathrm{\alpha/M}$] (right panel) obtained from APOGEE. The upper panels show the predictions versus the target and the black line indicates the identity line. The lower panels represent the residuals between the SIDRA-RVS’s $\mathrm{log}_{10}(\tau_{\mathrm{SIDRA}}$ $\mathrm{[Gyr]})$ prediction and BINGO’s true $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ denoted as $\mathrm{log}_{10}(\frac{\tau_{\mathrm{SIDRA}}}{\tau_{\mathrm{BINGO}}})$. The black filled circles and vertical error bars indicate the mean and the standard deviation of the residuals at the different $\mathrm{log}_{10}(\tau_{\mathrm{BINGO}}$ $\mathrm{[Gyr]})$ bins, respectively.
The colours of the dots in Figs. 2 and 3 indicate [M/H] (left panel) and $\mathrm{\alpha/M}$] (right panel). In general, the older age stars have lower [M/H] and higher $\alpha/\mathrm{M}]$, and the BINGO ages show the clear correlation with [M/H] and $\alpha/\mathrm{M}]$, as shown in Ciucă et al. (2021). The SIDRA ages also follow the trend. However, it is not as strong as the trend seen for the BINGO ages, and the stars with similar SIDRA age have a wider variety of [M/H] and $\alpha$/M]. This ensures that SIDRA is not solely learning the $\mathrm{\alpha/M}$]-age or [M/H]-age trends, as these trends are more pronounced for BINGO ages.